(10 pts) A police car siren emits 120W of sound at a frequency of 1200Hz a. What is the intensity of the sound in decibels at a distance of 500m from the siren? b. If the police car, travelling east at 200 km/hr chases a speeding motorist going east at 140 km/hr. What is the siren frequency as heard by the motorist?

8. A 4.00 mm think glass window has thermal conductivity k 1.00 W/(m K). What is the rate of heat loss through a 90.0 cm x 60.0 cm window by conduction if the indoor temperature is 60.0°F and the outdoor temperature is 20.0°F (2990 W)
Answer #1

7.

(a)

I=\frac {P}{4\pi r^2}=\frac {120}{4\pi \cdot 500^2}\: W/m^2= 3.82\times 10^{-5}\: W/m^2

(b)

f=\frac {V-v_{m}}{V+v_{source}}f_{o}

f=\frac {343-140}{343+200}\cdot 1200\: Hz= 448.62\: Hz

8.

R=\frac {KA(T_2-T_1)}{l}=\frac {1\cdot 90\cdot 60\cdot 10^{-4}\cdot (288.70-266.48)}{0.004}\: J/s\approx 2990\: W